OpenStudy (anonymous):

the factor 3m^3+5m^2-12m-20 help me plezz i have one day to finsh last resort

OpenStudy (cathyangs):

Ok, well I would start out by looking at the coefficients. Those are: 3, 5, -12, and -20. Do any of those have any factors in common? :)

OpenStudy (anonymous):

3 has -12in common and 5 has -20 in common

OpenStudy (cathyangs):

Yes, so if you group the ones that have something in common with each other together...

OpenStudy (anonymous):

what im lost

OpenStudy (cathyangs):

If you group the monomials (3m^3 is one monomial) that have factors in common together...you're basically just grouping the 3-coefficient monomial with the -12 coefficient monomial, and the same for the 5 and -20 ones. Then you can divide out the factor that each group has in common.

OpenStudy (anonymous):

so for 3and -12 the grouping would be -4 and the grouping for 5 and -20 is -4 right?

OpenStudy (cathyangs):

Yes! So if you use the "reverse distributive property" on each group?

OpenStudy (anonymous):

reverse distributive property ?

OpenStudy (cathyangs):

If you apply the distributive property, but reverse it. Instead of multiplying everything in parenthesis by the same thing, you divide "out" a common factor.

OpenStudy (anonymous):

so (3-20)4 like that

OpenStudy (cathyangs):

um...Not quite. Do you know what a factor is? The GCF, or Greatest Common Factor of two numbers is the largest number that both of those numbers can both be evenly divide by. To put it simply, here's an example: the GCF of 40 and 8 is 8, because both 40 and 8 are evenly divisible y 8, with no remainders. Now fiind the GCF for 3 and -12 and the numbers 5 and -20

OpenStudy (anonymous):

3 and 5

OpenStudy (cathyangs):

That's right. And now, can you group the numbers, and divide out the GCF for each group?

OpenStudy (anonymous):

so 3 in to 3 is one 3 in to 12 is 4 5 in to 5 is one 5 in to 20 is 4

OpenStudy (cathyangs):

Yes. Can you rewrite each group with parenthesis around the quotients? ( the 3 and 4 in the first group, for example)

OpenStudy (anonymous):

(3/3)=1 (3/12)=4 (5/5)=1 (5/20)=4

OpenStudy (cathyangs):

XD ok, that's one way to write it. Here's more of what I was looking for: For the first group: 3m^3-12m is rewritten as 3 (m^3-4) Try the second group.

OpenStudy (anonymous):

5m^2-20 5(m^2-4

OpenStudy (anonymous):

so was that right

OpenStudy (anonymous):

cathyangs u there

OpenStudy (cathyangs):

Sorry, just a family thing. That's right! (but remember to close the parenthesis) Now look back at the first one. What other factor do the two monomials inside the parenthesis have in common? (Big 'ol hint, it's a variable)

OpenStudy (anonymous):

m

OpenStudy (cathyangs):

And you can pull that out of the parenthesis, right? :) So now, write the two groups added together (since they were two parts of the same equation in the beginning)

OpenStudy (anonymous):

3(m^3-4)5(m^2-20)

OpenStudy (anonymous):

like this

OpenStudy (cathyangs):

*added together. you need a plus sign in there...also, double check each of the "groups" You forgot the m in the first one, and -20 should be something else in the second one ;)